Nemesis, Inc., has 210,000 shares of stock outstanding. Each share is worth $80, so the company’s market value of equity is $16,800,000. Suppose the firm issues 46,000 new shares at the following prices: $80, $74, and $68.

What will be the ex-rights price and the effect of each of these alternative offering prices on the existing price per share? (Leave no cells blank; if there is no effect select "No change" from the dropdown and enter "0". Round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

Current share price of Nemesis, Inc. = $80 per share

Number of share outstanding = 210,000 shares

Therefore, Market Value of shares prior to issue of new shares = $80 * 210,000 = $16,800,000

Number of share for new right issue = 46,000 shares

If Price of new issue = $80 per share

Cash raised from new issue = $80 * 46,000 = $3,680,000

Ex-rights stock price = (Market Value of shares prior to rights issue + Cash raised from new issue)/ Total Number of shares after the new issue

= ($16,800,000 + $3,680,000)/ (210,000 + 46,000)

= $20,480,000/ 256,000

= $80.00 per share

If Price of new issue = $74 per share

Cash raised from new issue = $74 * 46,000 = $3,404,000

Ex-rights stock price = (Market Value of shares prior to rights issue + Cash raised from new issue)/ Total Number of shares after the new issue

= ($16,800,000 + $3,404,000)/ (210,000 + 46,000)

= $20,204,000/ 256,000

= $78.92 per share

If Price of new issue = $68 per share

Cash raised from new issue = $68 * 46,000 = $3,128,000

Ex-rights stock price = (Market Value of shares prior to rights issue + Cash raised from new issue)/ Total Number of shares after the new issue

= ($16,800,000 + $3,128,000)/ (210,000 + 46,000)

= $19,928,000/ 256,000

= $77.84 per share